OpenStudy (anonymous):

Write the equations in standard form (I'm working with parabolas): x=-1/8 y^2 x-5=4(y-3)^2 x^2-2y-6x=-5 y=-x^2

OpenStudy (blurbendy):

do you know what the standard form for a parabola is?

OpenStudy (anonymous):

In my notes I have different equations for horizontal and vertical parabolas but I don't know what the standard form for the parabola is. @blurbendy

OpenStudy (blurbendy):

standard form for a parabola is y = ax^2 + bx + c

OpenStudy (anonymous):

Okay. I'm sure I can figure out how to put the equations into that form, but how do I figure out the focal point and directrix from that equation?

OpenStudy (anonymous):

@blurbendy

OpenStudy (blurbendy):

it's been a while since I've done these, so I don't quite remember. I should also mention that if a parabola is "sideways" then the equation is x = ay^2 + by + c

OpenStudy (anonymous):

Okay, thank you

OpenStudy (blurbendy):

sorry, guess I need to retake algebra 2 XD

OpenStudy (blurbendy):

actually, I just looked it up. it's not that hard. let's start with the last one. y = -x^2 form: x^2 = (4p)y y = -p (directrix) focus: (0, p) step 1: get -x^2 by itself -y = x^2 step 2: 4p = -1 p = -1/4 y = -p, y = 1/4, directrix = 1/4 focus = (0, -1/4)

OpenStudy (blurbendy):

let me know if you have any questions

OpenStudy (anonymous):

@blurbendy For the 3rd equation how would I get it into a form so I can figure out what p is?

OpenStudy (blurbendy):

youll want to get it in y = ax^2 + bx + c form

OpenStudy (blurbendy):

get y by itself

OpenStudy (anonymous):

so it would be y = 1/2(x^2 - 6x +5) ? and then to find p it would be 4p = 1/2 ?

OpenStudy (blurbendy):

the standard form looks correct, but Im not sure how to find p